OpenStudy (anonymous):

Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received. x - 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive?

OpenStudy (anonymous):

93750

OpenStudy (anonymous):

\[x-62,500=0.25y\]Simply substitute y with 125,000\[x-62,500=0.25\times125,000\]\[x-62,500=31,250\]\[x=93,750\]So, Maya received 93,750 votes.